2 point charges of 20C and80C are placed 18cm apart.Findd the position of the point where E^ is 0

Grade:12

1 Answers

Vikas TU
7 years ago
E would be zero b/w both the chrges as both are psotive and would exhibits the repelling forces.
Hence, let at x distance the Electric field would be  due to both =>
E net = E1 + E2 = 0
20k/x^2 = 80k/(18 -x)^2
18 – x = 2x
3x = 18
x = 6 cm.
Hence at 6 cm from 20 C charge ot would be zero.